OpenStudy (anonymous):

Given this equation y= 2 x^2- 4 x-3. Complete the square to put it in standard form, i.e. in the form y=a(x-h)^2+k.

OpenStudy (anonymous):

y= 2 x^2- 4 x-3 y=2(x^2-2x +1 -1)-3 y=2(x^2-2x +1)-2-3 y=2(x-1)^2 + (-5)